What is the number of electron in the piece of iron with a mass of 0.70g
1
Expert's answer
2020-04-13T02:34:44-0400

Iron is 26th element in Periodic table. It means that atoms of this element have 26 protons in nucleus and therefore 26 electrons in each atom.

m(Fe) = 0.70 g

n(Fe) = 0.70/56 = 0.0125 mol

number of atoms N(Fe) = n(Fe)*NA = 0.0125*6.02*1023 = 7.525*1021

number of electrons in this amount of atoms is N(e) = N(Fe)*26 = 7.525*1021*26 =1.9565*1023